OpenStudy (anonymous):

How many odd natural numbers of three digits are possible if repetition of digits is allowed? (Use combination)

OpenStudy (gudden):

so r Prime numbers also allowed @CO_oLBoY

OpenStudy (anonymous):

yup all odd natural num are

OpenStudy (anonymous):

i think its like\[\left(\begin{matrix}3 \\ 3\end{matrix}\right) + \left(\begin{matrix}5 \\ 3\end{matrix}\right) + ........ + \left(\begin{matrix}2n - 1 \\ 3\end{matrix}\right)\] Am I correct??

OpenStudy (gudden):

three digit numbers start from 100 and end with 999 from 100 - 200 number of odd numbers = 50 from 200 - 300 , number of odd numbers = 50 from 300 - 400 , number of odd number = 50 from 400 - 500 , number of odd numbers = 50 from 500 - 600 , number of odd numbers = 50 from 600 - 700 , number of odd numbers = 50 from 700 - 800 , no of odd no's = 50 from 800 - 900, no of od no's = 50 and from 900 - 999 , no of odd number's = 50 ( so can u add those now) i guess, u could use this one too :p

OpenStudy (anonymous):

@gudden but we have to use combinition

OpenStudy (gudden):

u didn't mention it... @CO_oLBoY

OpenStudy (anonymous):

sorry

OpenStudy (gudden):

its alright..

OpenStudy (anonymous):

@gudden cz they dont mention what method u should use. so u've got to find out urself

OpenStudy (***[isuru]***):

lol... i ve got this SLMC by april 15 and the paper include combinatories... at the moment im trying to understand this subject at my top speed from a book with 645 pgs... but im only 50 pgs up ... so im not much of a use here :(

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

no prblm

OpenStudy (anonymous):

so @gudden wht do u think the sol. is?

OpenStudy (gudden):

idk something about combinations... i have a few books realted with quantitative and qualitative reasoning...and would try to find about it in those.. so, wont u mind if i could give the answer tomorrow.. (bt am not sure , if i would get it in any of those books or not)

OpenStudy (anonymous):

take ur time

OpenStudy (gudden):

my time would be around 10 hours and still no surity if i would get it or not..

OpenStudy (gudden):

try it http://www.algebra.com/algebra/homework/Permutations/

OpenStudy (anonymous):

@ganeshie8 what do u think?

ganeshie8 (ganeshie8):

odd numbers end in 1, 3, 5, 7, 9

ganeshie8 (ganeshie8):

5 choices for last digit 10 choices for middle digit 9 choices for first digit

OpenStudy (anonymous):

is my ans at the start right?

OpenStudy (anonymous):

is this right? \[\left(\begin{matrix}3 \\ 3\end{matrix}\right) + \left(\begin{matrix}5 \\ 3\end{matrix}\right) + ....... + \left(\begin{matrix}2n - 1 \\ 3\end{matrix}\right)\]

OpenStudy (anonymous):

@ParthKohli got any solution?

Parth (parthkohli):

ganeshie is right.

OpenStudy (anonymous):

i think i've also mentioned we have to use combination man

OpenStudy (kainui):

Honestly you don't have to use combination for this one. Just as long as you realize that if you count to 1000 you had to go through 0 and 99 then you have 1000 so you can subtract off those numbers there. Then odds happen every other, so divide by 2. Just a quick sanity check of whatever answer you do end up with.

OpenStudy (kainui):

Since you're allowing repeats, the factorial function won't help you out @CO_oLBoY

OpenStudy (anonymous):

its the question that demands to use combination

OpenStudy (kainui):

I think you're imposing a false condition on yourself based on what you think "combination" means.

OpenStudy (anonymous):

just tell me the ans maaaaaaaaaaaaaaaaaan is mine right???

OpenStudy (kainui):

If I say I have three colors of balls, Red, Blue, and Green and that I also have 3 spaces to place balls, then if I DON'T use the same color multiple times I can place them 3! ways since the first spot I have 3 choices, then the second has 2, the last has 1. But if I am allowed to REPEAT colors then I have 3 choices for the first, 3 choices for the second, and 3 choices for the last. Completely different. Both scenarios are combinations.

OpenStudy (kainui):

I don't even understand what your answer is.

OpenStudy (anonymous):

when repition is allowed its combination. i it aint then its permutation @Kainui